# Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback

### v1

@@ Abstract
Figure 3: The behavior of ${\dot{W}_{\mathrm{ext},u}}/{T_{u}}+{\dot{W}_{\mathrm{ext},v}}/{T_{v}}$ as a function of the feedback gain $g$ in the steady state of Eq. (Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback). The feedback protocol is based on the quantum Kalman filter, where the feedback Hamiltonian is given by $H^{\hat{x},\hat{p}}_{\rm{FB}}=\hbar g(-\langle x\rangle_{\rm{c}}\hat{p}+\langle p\rangle_{\rm{c}}\hat{x})$ (orange curve). In addition, we demonstrate another feedback Hamiltonian $H^{\hat{x}}_{\rm{FB}}=\hbar g(\langle x\rangle_{\rm{c}}+\langle p\rangle_{\rm{c}})\hat{x}$ (blue dashed curve) for the sake of comparison. The behavior of ${\dot{W}_{\rm{ext}}}/{T}$ is also shown (green curve). The black line represents the QC-transfer entropy, the bound of  (8), to which only the orange curve approaches in large $g$. The parameters are chosen as $k/\omega=0.18$, $\eta=0.34$, $T=292$K, and $\bar{n}\gamma/\omega=0.0058$ Magrini et al. (2021). The vertical axis is normalized by $\dot{I}_{\mathrm{QCT}}$
Numerical demonstration.— We present a numerical demonstration of our main inequality (8), by numerically implementing the quantum Kalman filter. Figure 3 shows the behavior of ${\dot{W}_{\mathrm{ext},u}}/{T_{u}}+{\dot{W}_{\mathrm{ext},v}}/{T_{v}}$ and ${\dot{W}_{\rm{ext}}}/{T}$ as functions of the feedback gain in the steady state of Eq. (Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback). The parameters are taken from Ref. Magrini et al. (2021), an experiment of feedback cooling of levitated nanoparticles.
Figure 3 compares the two feedback Hamiltonians and shows that only $H^{\hat{x},\hat{p}}_{\rm{FB}}=\hbar g(-\langle x\rangle_{\rm{c}}\hat{p}+\langle p\rangle_{\rm{c}}\hat{x})$, plotted in the orange curve, ensures that ${\dot{W}_{\mathrm{ext},u}}/{T_{u}}+{\dot{W}_{\mathrm{ext},v}}/{T_{v}}$ approaches the upper bound $\dot{I}_{\mathrm{QCT}}$, as $g\to\infty$. Since feedback proportional to $\hat{x}$ is commonly used in experiments, we also consider the Hamiltonian of the form $H^{\hat{x}}_{\rm{FB}}=\hbar g(\langle x\rangle_{\rm{c}}+\langle p\rangle_{\rm{c}})\hat{x}$ (blue dashed curve), but it does not allow the system to reach the information-theoretic bound of (8). Moreover, $\dot{W}_{\rm{ext}}/T$ with feedback Hamiltonian $H^{\hat{x},\hat{p}}_{\rm{FB}}$ (green curve) fails to asymptotically approach $\dot{I}_{\rm QCT}$, indicating that the naive bound without using the kinetic temperatures is not achievable.
Conclusion and outlook.— In this study, we established a fundamental limit of quantum feedback cooling in Gaussian systems under continuous QND measurement, by linking work extraction, kinetic temperatures, and QC-transfer entropy (8). Since the derived inequality is based on the kinetic temperatures rather than the bath temperature, it imposes a stricter constraint on cooling performance than known information-thermodynamic inequalities, and can be saturated through a sufficiently large feedback gain with the quantum Kalman filter (Fig. 3). In addition, we discussed the relationship between QC-transfer entropy and yet another quantity called QC information flow, as characterized by inequality (12). These results elucidate to what extent information acquired from the measurement can enhance the extraction of the work and reduce the kinetic temperatures.
Our results can be experimentally verified using current or near-future experimental technologies. For example, in levitated nanoparticle systems, feedback cooling based on the quantum Kalman filter has been realized Magrini et al. (2021), making experimental verification promising. However, current experiments typically utilize feedback proportional only to the position operator, whereas achieving the theoretical limit requires feedback proportional to the momentum operator as well. Realizing this optimal feedback is experimentally challenging, but is anticipated to lead to the verification of the fundamental connection between information and thermodynamics in feedback cooling.
Acknowledgements.
Acknowledgments.— We are grateful to Kiyotaka Aikawa and Isaac Layton for valuable discussions. This work was supported by JST ERATO Grant No. JPMJER2302, Japan. K.K. and T.Y. are supported by World-leading Innovative Graduate Study Program for Materials Research, Information, and Technology (MERIT-WINGS) of the University of Tokyo. T.Y. is also supported by JSPS KAKENHI Grant No. JP23KJ0672. K.F. is supported by JSPS KAKENHI Grant Nos. JP23K13036 and JP24H00831. T.S. is supported by JSPS KAKENHI Grant No. JP19H05796 and JST CREST Grant No. JPMJCR20C1. T.S. is also supported by Institute of AI and Beyond of the University of Tokyo.
Supplemental Material for “Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback”
Kousuke Kumasaki1, Toshihiro Yada1, Ken Funo1, and Takahiro Sagawa1,2
1Department of Applied Physics, The University of Tokyo, 7-3-1 Hongo, Bunkyo-ku, Tokyo 113-8656, Japan
2Quantum-Phase Electronics Center (QPEC), The University of Tokyo, 7-3-1 Hongo, Bunkyo-ku, Tokyo 113-8656, Japan
I Quantum information flow in general setup
In this section, we introduce the measures of quantum information flow and derive the generalized second law, in the setup of multiple quantum measurement and feedback, which includes the continuous feedback case, mainly addressed in this manuscript, as a special case. In Sec. I.1, we introduce QC information flow, and derive the generalized second law. In Sec. I.2, we introduce QC-transfer entropy and discuss its relationship with QC information flow. In Sec. I.3, we demonstrate that the continuous feedback control setup discussed in the main text is a special case of the general multiple feedback setup.
I.1 Generalized second law with QC information flow
Figure S1: Schematic for the multiple measurement and feedback setup. We consider the situation where the measurement and feedback are alternately performed. The $n$-th measurement outcome is represented as $y_{t_{n}}$, and all the measurement outcomes until $y_{t_{n}}$ is denoted as $Y_{t_{n}}\coloneqq(y_{t_{0}},y_{t_{1}},\dots,y_{t_{n}})$. The state right before and after the $n$-th measurement are denoted as $\hat{\tau}_{t_{n}}^{Y_{t_{n}}}$ and $\hat{\rho}_{t_{n+1}}^{Y_{t_{n+1}}}$, respectively, whose superscript represents the corresponding measurement outcomes. The superoperator for $n$-th feedback process is described as $\mathcal{E}_{t_{n}}^{e_{t_{n}}}$, where $e_{t_{n}}$ represents the estimator determined by $Y_{t_{n}}$.
We here consider the general setup of multiple measurement and feedback, as shown in Fig. S1. The process starts at the time $t=t_{0}$, and the $n$-th measurement and feedback is conducted during the time interval $[t_{n},t_{n+1})$. The $n$-th measurement outcome is represented as $y_{t_{n}}$, and all the measurement outcomes until $y_{t_{n}}$ is denoted as $Y_{t_{n}}\coloneqq(y_{t_{0}},y_{t_{1}},\dots,y_{t_{n}})$. The conditional density operators right before and after the $n$-th measurement are denoted as $\hat{\tau}_{t_{n}}^{Y_{t_{n}}}$ and $\hat{\rho}_{t_{n+1}}^{Y_{t_{n+1}}}$, respectively, whose superscript represents the corresponding measurement outcomes. The $n$-th feedback process depends on the $n$-th estimator, which is determined as a function of $Y_{t_{n}}$ as $e_{t_{n}}\coloneqq h(Y_{t_{n}})$, and $\mathcal{E}_{t_{n}}^{e_{t_{n}}}$ is the superoperator representing the feedback process. The density operator conditioned on the estimator $e_{t_{n}}$ is represented as $\hat{\rho}_{t_{n}}^{e_{t_{n}}}$, where the relationship $\sum_{h(Y_{t_{n}})=e_{t_{n}}}P(Y_{t_{n}})\hat{\rho}_{t_{n}}^{Y_{t_{n}}}=P(e_{t_{n}})\hat{\rho}_{t_{n}}^{e_{t_{n}}}$ is satisfied. The feedback process consists of the heat-bath interaction and external driving, and therefore the superoperator can be decomposed as $\mathcal{E}_{t_{n}}^{e_{t_{n}}}=\mathcal{E}^{e_{t_{n}}}_{\rm drv}\circ\mathcal{E}^{e_{t_{n}}}_{\rm bath}$.
$I_{\mathrm{QCI}}^{S}\coloneqq\sum_{n=n_{i}}^{n_{f}-1}\chi\left(\hat{\rho}_{t_{n+1}}:e_{t_{n}}\right)-\chi\left(\hat{\rho}_{t_{n}}:e_{t_{n}}\right),$ (S1)
where $\chi(\hat{\rho}:e)\coloneqq S(\hat{\rho})-\mathbb{E}_{e}[S(\hat{\rho}^{e})]$ represents the Holevo information Holevo (1973), $S(\hat{\rho})\coloneqq-\mathrm{tr}[\hat{\rho}\ln\hat{\rho}]$ is the von Neumann entropy, $\mathbb{E}_{X}$ represents taking the expectation value over a random variable $X$, and $\hat{\rho}_{t_{n}}\coloneqq\mathbb{E}_{e_{t_{n}}}[\hat{\rho}_{t_{n}}^{e_{t_{n}}}]$. Since the Holevo information quantifies the correlation between the quantum system and the (classical) estimator, $-I_{\mathrm{QCI}}^{S}$ is the total information utilized in the multiple feedback. The effect of measurement backaction is quantified as
$S_{\rm BA}\coloneqq\sum_{n=n_{i}}^{n_{f}-1}\mathbb{E}_{e_{t_{n}}}\left[S\left(\hat{\rho}_{t_{n+1}}^{e_{t_{n}}}\right)-S\left(\hat{\tau}_{t_{n}}^{e_{t_{n}}}\right)\right].$ (S2)
$Q\coloneqq\sum_{n=n_{i}}^{n_{f}-1}\mathbb{E}_{e_{t_{n}}}[Q^{e_{t_{n}}}]=\sum_{n=n_{i}}^{n_{f}-1}\mathbb{E}_{e_{t_{n}}}\left[\mathrm{tr}\left[H(\mathcal{E}^{e_{t_{n}}}_{\rm bath}(\hat{\rho}^{e_{t_{n}}}_{t_{n}})-\hat{\rho}^{e_{t_{n}}}_{t_{n}})\right]\right],$ (S3)
$\begin{split}\beta(W_{\rm ext}+\Delta F)+I_{\mathrm{QCI}}^{S}+S_{\rm BA}&=-S(\hat{\rho}_{t_{n_{f}}})+S(\hat{\rho}_{t_{n_{i}}})+\beta Q+I_{\mathrm{QCI}}^{S}+S_{\rm BA}\\ &=-\sum_{n={n_{i}}}^{n_{f}-1}\left\{\mathbb{E}_{e_{t_{n}}}\left[S\left(\hat{\tau}_{t_{n}}^{e_{t_{n}}}\right)-S\left(\hat{\rho}_{t_{n}}^{e_{t_{n}}}\right)-\beta Q^{e_{t_{n}}}\right]\right\}\\ &=-\sum_{n={n_{i}}}^{n_{f}-1}\left\{\mathbb{E}_{e_{t_{n}}}\left[S\left(\mathcal{E}^{e_{t_{n}}}_{\rm bath}(\hat{\rho}^{e_{t_{n}}}_{t_{n}})\right)-S\left(\hat{\rho}_{t_{n}}^{e_{t_{n}}}\right)-\mathrm{tr}\left[\left(\mathcal{E}^{e_{t_{n}}}_{\rm bath}(\hat{\rho}^{e_{t_{n}}}_{t_{n}})-\hat{\rho}^{e_{t_{n}}}_{t_{n}}\right)\ln\hat{\rho}^{\mathrm{G}}\right]\right]\right\}\\ &=-\sum_{n={n_{i}}}^{n_{f}-1}\left\{\mathbb{E}_{e_{t_{n}}}\left[S(\hat{\rho}^{e_{t_{n}}}_{t_{n}}\|\hat{\rho}^{\mathrm{G}})-S(\mathcal{E}^{e_{t_{n}}}_{\rm bath}(\hat{\rho}^{e_{t_{n}}}_{t_{n}})\|\hat{\rho}^{\mathrm{G}})\right]\right\}\\ &\leq 0,\end{split}$ (S4)
where $S(\hat{\rho}\|\hat{\sigma})\coloneqq\mathrm{tr}[\hat{\rho}(\ln\hat{\rho}-\ln\hat{\sigma})]$ represents the quantum relative entropy, and $\hat{\rho}^{\mathrm{G}}\propto\exp(-\beta H)$ is a Gibbs state with the inverse temperature $\beta$. The equality in the third line follows from the fact that $\hat{\tau}_{t_{n}}^{e_{t_{n}}}=\mathcal{E}^{e_{t_{n}}}_{\rm drv}\circ\mathcal{E}^{e_{t_{n}}}_{\rm bath}(\hat{\rho}^{e_{t_{n}}}_{t_{n}})$ with the superoperator $\mathcal{E}^{e_{t_{n}}}_{\rm drv}$ representing purely unitary driving, and therefore $S(\hat{\tau}_{t_{n}}^{e_{t_{n}}})=S(\mathcal{E}^{e_{t_{n}}}_{\rm bath}(\hat{\rho}^{e_{t_{n}}}_{t_{n}}))$. The inequality in the final line follows from the fact that $\hat{\rho}^{\mathrm{G}}$ is the fixed point of the heat-bath interaction $\mathcal{E}^{e_{t_{n}}}_{\rm bath}$, and the monotonicity of quantum relative entropy under a CPTP (completely positive and trace-preserving) map Petz (2003).
I.2 QC-transfer entropy
$I_{\mathrm{QCT}}\coloneqq\sum_{n={n_{i}}}^{n_{f}-1}\mathbb{E}_{Y_{t_{n}}}\left[I_{\mathrm{QC}}\left(\hat{\tau}_{t_{n}}^{Y_{t_{n}}}:y_{t_{n+1}}\right)\right],$ (S5)
$I_{\mathrm{QCT}}-\Delta\chi\geq-I_{\mathrm{QCI}}^{\mathrm{S}}-S_{\mathrm{BA}},$ (S6)
$\begin{split}I_{\mathrm{QCT}}-\Delta\chi+I_{\mathrm{QCI}}^{\mathrm{S}}+S_{\rm BA}&=\sum_{n=n_{i}}^{n_{f}-1}\mathbb{E}_{Y_{t_{n}}}\left[S\left(\hat{\rho}_{t_{n}}^{Y_{t_{n}}}\|\hat{\rho}_{t_{n}}^{e_{t_{n}}}\right)-S\left(\hat{\tau}_{t_{n}}^{Y_{t_{n}}}\|\hat{\tau}_{t_{n}}^{e_{t_{n}}}\right)\right]\\ &\geq 0,\end{split}$ (S7)
where the first line follows from a straightforward rearrangement of terms, and the second line follows from the equalities $\mathcal{E}_{t_{n}}^{e_{t_{n}}}(\hat{\rho}_{t_{n}}^{e_{t_{n}}})=\hat{\tau}_{t_{n}}^{e_{t_{n}}}$ and $\mathcal{E}_{t_{n}}^{e_{t_{n}}}(\hat{\rho}_{t_{n}}^{Y_{t_{n}}})=\hat{\tau}_{t_{n}}^{Y_{t_{n}}}$, along with the monotonicity of quantum relative entropy under a CPTP map. As can be seen from Eq. (S7), a simple equality condition for Eq. (S6) is that $\hat{\rho}_{t_{n}}^{e_{t_{n}}}=\hat{\rho}_{t_{n}}^{Y_{t_{n}}}$ holds for all $Y_{t_{n}}$. This condition implies that the feedback estimator $e_{t_{n}}=h(Y_{t_{n}})$ is optimally designed, which is known to be achieved by a Kalman filter.
I.3 Correspondence to the continuous feedback setup considered in the main text
We note that the definitions of thermodynamic quantities and information terms in the main text are consistent with those in Secs. I.1 and I.2. In particular, the consistency of the heat income $Q$, work extraction $W_{\rm ext}$, and QC information flow $I_{\mathrm{QCI}}^{\rm S}$ can be directly verified by substituting the discrete time index $t_{n}$ with the continuous time variable $t$ as is demonstrated in Sec. II.2. While the correspondence is less straightforward for the measurement backaction term $S_{\mathrm{BA}}$ [Eq. (11) in the main text] and the QC-transfer entropy $I_{\mathrm{QCT}}$ [Eq. (4) in the main text], it can still be demonstrated by neglecting terms of order $O(dt)$. It is comprehensive to represent Eqs. (S2) and (S5) in the following forms taking the limit of $dt\to 0$:
$\dot{S}_{\mathrm{BA}}$ $=\mathbb{E}_{e_{t}}\left[\left.\lim_{dt\to 0}\frac{S(\mathcal{M}_{\mathrm{BA},t}(\hat{\rho}))-S(\hat{\rho})}{dt}\right|_{\hat{\rho}=\hat{\tau}_{t}^{e_{t}}}\right],$ (S8)
$\dot{I}_{\mathrm{QCT}}$ $=\mathbb{E}_{Y_{t}}\left[\left.\lim_{dt\to 0}\frac{I_{\rm QC}(\hat{\rho}:dy(t))}{dt}\right|_{\hat{\rho}=\hat{\tau}_{t}^{Y_{t}}}\right].$ (S9)
From the equalities $\hat{\tau}_{t}^{e_{t}}=\hat{\rho}_{t}^{e_{t}}-\frac{i}{\hbar}\left[\hat{H}_{\mathrm{HO}}+\hat{H}_{\mathrm{FB}}(t),\hat{\rho}_{t}^{e_{t}}\right]dt+\mathcal{L}_{\mathrm{bath}}[\hat{\rho}_{t}^{e_{t}}]dt$ and $\hat{\tau}_{t}^{Y_{t}}=\hat{\rho}_{t}^{Y_{t}}-\frac{i}{\hbar}\left[\hat{H}_{\mathrm{HO}}+\hat{H}_{\mathrm{FB}}(t),\hat{\rho}_{t}^{Y_{t}}\right]dt+\mathcal{L}_{\mathrm{bath}}[\hat{\rho}_{t}^{Y_{t}}]dt,$ and the fact that the terms in the square brackets are smooth at the respective points $\hat{\rho}=\hat{\tau}_{t}^{e_{t}}$ and $\hat{\rho}=\hat{\tau}_{t}^{Y_{t}}$, we can replace $\hat{\tau}_{t}^{e_{t}}$ and $\hat{\tau}_{t}^{Y_{t}}$ in these equations with $\hat{\rho}_{t}^{e_{t}}$ and $\hat{\rho}_{t}^{Y_{t}}$, respectively, introducing only negligible additional terms of order $O(dt)$. We also note that in the main text, we use the notations $\hat{\rho}_{\rm c}(t)=\hat{\rho}_{t}^{Y_{t}}$ and $\hat{\rho}_{t,\mathrm{BA}}^{e_{t}}=\mathcal{M}_{\mathrm{BA},t}(\hat{\rho}_{t}^{e_{t}})$ for simplicity. In Sec. II, we use the expressions (S8) and (S9) to derive the main results presented in the main text.
II Information thermodynamic cooling limit in Gaussian systems
In this section, we present the derivation of inequality (10) in the main text:
$\frac{\dot{W}_{\mathrm{ext},u}}{T_{u}}+\frac{\dot{W}_{\mathrm{ext},v}}{T_{v}}\leq-\dot{I}_{\mathrm{QCI}}^{S}-\dot{S}_{\rm BA}.$ (S10)
We further show the following relation (12) in the main text under the steady-state condition:
$-\dot{I}_{\mathrm{QCI}}^{S}-\dot{S}_{\rm BA}\leq\dot{I}_{\rm QCT}.$ (S11)
Inequalities (10) and (S11) are derived for a Gaussian system that obeys the following stochastic master equation:
$d\hat{\rho}_{t}^{Y_{t}}=-\frac{i}{\hbar}\left[\hat{H}_{\mathrm{HO}}+\hat{H}_{\mathrm{FB}}(t),\hat{\rho}_{t}^{Y_{t}}\right]dt+\mathcal{L}_{\mathrm{bath}}[\hat{\rho}_{t}^{Y_{t}}]dt+2k\mathcal{D}_{\hat{x}}[\hat{\rho}_{t}^{Y_{t}}]dt+\sqrt{2k\eta}\mathcal{H}_{\hat{x}}[\hat{\rho}_{t}^{Y_{t}}]dW.$ (S12)
We note that the dynamics of the system described by Eq. (Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback) is a special case of the general setup discussed in Sec. I.1, where the explicit form of the superoperators for the heat-bath interaction, external driving, and measurement backaction are respectively given by:
$\mathcal{E}^{e_{t}}_{\rm bath}(\hat{\rho})$ $=\hat{\rho}-\frac{i}{\hbar}\left[\hat{H}_{\mathrm{HO}},\hat{\rho}\right]dt+\mathcal{L}_{\mathrm{bath}}[\hat{\rho}]dt,$ (S13)
$\mathcal{E}^{e_{t}}_{\rm drv}(\hat{\rho})$ $=\hat{\rho}-\frac{i}{\hbar}\left[\hat{H}_{\mathrm{FB}}(t),\hat{\rho}\right]dt,$ (S14)
$\mathcal{M}_{\mathrm{BA},t}(\hat{\rho})$ $=\hat{\rho}+2k\mathcal{D}_{\hat{x}}[\hat{\rho}]dt,$ (S15)
$\hat{\rho}_{t}^{e_{t}}\rightarrow\mathcal{E}^{e_{t}}_{\rm bath}(\hat{\rho}_{t}^{e_{t}})=:\hat{\pi}_{t}^{e_{t}}\rightarrow\mathcal{E}^{e_{t}}_{\rm drv}(\hat{\pi}_{t}^{e_{t}})=:\hat{\tau}_{t}^{e_{t}}\rightarrow\mathcal{M}_{\mathrm{BA},t}(\hat{\tau}_{t}^{e_{t}}):=\hat{\rho}_{t+dt}^{e_{t}}.$ (S16)
We note that the Hamiltonian of the system $H$ is given by $\hat{H}_{\rm HO}=(\hbar\omega/2)(\hat{x}^{2}+\hat{p}^{2})$ and the feedback Hamiltonian is defined as $\hat{H}_{\mathrm{FB}}(t)=\hbar c_{x}(e_{t})\hat{x}-\hbar c_{p}(e_{t})\hat{p}$, where $c_{x}$ and $c_{p}$ are linear functions of estimate $e_{t}$, defining the feedback coefficients. We further assume that the feedback Hamiltonian is chosen so that, when the initial state satisfies $\langle\hat{x}\rangle(0)=\langle\hat{p}\rangle(0)=0$, the expectation values at later times satisfy $\langle\hat{x}\rangle(t)=\langle\hat{p}\rangle(t)=0$. Intuitively, such condition is satisfied when the feedback Hamiltonian is designed to reduce the mean position and momentum to the origin of the phase space. On the other hand, such condition is not satisfied when the feedback Hamiltonian realizes a constant displacement of the position $\hat{H}_{\mathrm{FB}}(t)=a\hat{x}$, where $a$ is a constant.
In what follows, we first derive the representation of kinetic temperatures using the Gaussian property of the quantum state in Sec. II.1, and then derive (10) in Sec. II.2. We further assume the steady-state condition and derive (S11) in Sec. II.3.
II.1 Kinetic temperatures
In what follows, we introduce the kinetic temperatures of one-mode Gaussian state $T_{u}$ and $T_{v}$. We start by noting that any single-mode Gaussian states centered at the origin can be regarded as a squeezed thermal state Kim et al. (1989) obtained by applying a squeezing operation $\hat{S}(\xi)$ to the Gibbs state $\hat{\rho}^{\mathrm{G}}$. In other words, for any single-mode Gaussian state $\hat{\rho}_{t}$ centered at the origin, there exists a pair $(\hat{\rho}^{\mathrm{G}},\xi)$ such that Olivares (2012):
$\hat{\rho}_{t}=\hat{S}(\xi)\hat{\rho}^{\mathrm{G}}\hat{S}(\xi)^{\dagger}.$ (S17)
Here, the squeezing operator is parameterized by the complex number $\xi=r\mathrm{e}^{i\theta}$ as
$\hat{S}(\xi):=\exp\left(\frac{1}{2}\xi(\hat{a}^{\dagger})^{2}-\frac{1}{2}\xi^{*}\hat{a}^{2}\right),$ (S18)
and the Gibbs state is expressed in terms of the inverse temperature $\beta^{\mathrm{G}}$ as
$\hat{\rho}^{\mathrm{G}}:=\mathrm{e}^{-\beta^{\mathrm{G}}\hat{H}_{\mathrm{HO}}}/Z(\beta^{\mathrm{G}}),$ (S19)
where $Z(\beta)$ is the partition function given by $Z(\beta^{\mathrm{G}}):=\mathrm{tr}[\mathrm{e}^{-\beta^{\mathrm{G}}\hat{H}_{\mathrm{HO}}}]$. Now, the explicit expression of the covariance matrix of the Gaussian state reads
$\sigma_{xx}$ $=[\cosh{(2r)}+\sinh{(2r)}\cos\theta]\sigma^{\rm G},$
$\sigma_{pp}$ $=[\cosh{(2r)}-\sinh{(2r)}\cos\theta]\sigma^{\rm G},$
$\sigma_{xp}$ $=\sinh{(2r)}\sin\theta\sigma^{\rm G},$ (S20)

### v2

@@ III.3 Outline of the proof, and more on quantum information flow
IV Achievement of the cooling limit based on the quantum Kalman filter
An example of experimental systems that obey the stochastic master equation (II.1) is the levitated nanoparticle system Magrini et al. (2021). In nanoparticle systems, the degree of cooling is often evaluated by how small the occupation number of the particle, $\langle n\rangle=\langle a^{\dagger}a\rangle$, can be made. Using the variances of the state, this can be expressed as $\braket{n}=\frac{1}{2}\left(\sigma_{xx}+\sigma_{pp}\right)-1/2$ . Defining the occupation number corresponding to $\bm{\sigma}_{\mathrm{c}}$ as $\langle n\rangle_{\min}\coloneqq\frac{1}{2}(\sigma_{xx,\mathrm{c}}+\sigma_{pp,\mathrm{c}})$, we obtain:
$\braket{n}\geq\braket{n}_{\mathrm{min}}:=\frac{1}{2}\left(\sigma_{xx,\mathrm{c}}^{2}+\sigma_{pp,\mathrm{c}}^{2}\right)-\frac{1}{2}.$ (21)
Here, $\langle n\rangle_{\min}$ represents the cooling limit. The equation of the inequality (21) is fulfilled when $\bm{\sigma}=\bm{\sigma}_{\mathrm{c}}$ holds. In the absence of a heat bath, the equality condition for (21) is known to be satisfied in the limit of strong feedback using the quantum Kalman filter estimates Doherty and Jacobs (1999). A similar sufficient condition can also be demonstrated in this case. At the same time, the equality condition of the inequality (14) holds. This means that the information thermodynamic bound obtained in this study and the cooling limit of nanoparticle systems are reached simultaneously. In what follows, we show the equality conditions of Eqs. (21) and (14).
We now briefly introduce the quantum Kalman filter, in which the estimated value $e_{t}=\{\braket{\hat{x}}_{\mathrm{c}},\braket{\hat{p}}_{\mathrm{c}}\}$ is obtained. We assume that a feedback Hamiltonian $\hat{H}_{\mathrm{FB}}(t)=-\left(a_{x}\braket{x}_{\rm{c}}+a_{p}\braket{p}_{\rm{c}}\right)\hat{p}+\left(b_{x}\braket{x}_{\rm{c}}+b_{p}\braket{p}_{\rm{c}}\right)\hat{x}$ is applied using this estimated value. Then, time evolutions of the expected values $\braket{\hat{x}}_{\mathrm{c}}$ and $\braket{\hat{p}}_{\mathrm{c}}$ governed by Eq. (II.1) are expressed as:
$d\braket{\hat{x}}_{\mathrm{c}}=$ $\left({{\omega}}\braket{\hat{p}}_{\mathrm{c}}-a_{x}\braket{\hat{x}}_{\mathrm{c}}-a_{p}\braket{\hat{p}}_{\mathrm{c}}-\frac{\gamma}{2}\braket{\hat{x}}_{\mathrm{c}}\right)dt$
$+8k\eta\sigma_{xx,\mathrm{c}}\cdot(dy-\braket{x}_{\mathrm{c}}dt),$ (22)
$d\braket{\hat{p}}_{\mathrm{c}}=$ $\left(-\omega\braket{\hat{x}}_{\mathrm{c}}-b_{x}\braket{\hat{x}}_{\mathrm{c}}-b_{p}\braket{\hat{p}}_{\mathrm{c}}-\frac{\gamma}{2}\braket{\hat{p}}_{\mathrm{c}}\right)dt$
$+8k\eta\sigma_{xp,\mathrm{c}}\cdot(dy-\braket{x}_{\mathrm{c}}dt).$ (23)
This stochastic differential equation allows us to determine the expected values of the position and momentum at the next time step based on current expected values and the measurement outcome $dy$. In the quantum Kalman filter, this mechanism is used to sequentially obtain the estimate $e_{t}=\{\braket{\hat{x}}_{\mathrm{c}},\braket{\hat{p}}_{\mathrm{c}}\}$. Here, the variance and covariance of the posterior state appearing in the filtering coefficients can be determined in advance because the differential equations of $\bm{\sigma}_{\mathrm{c}}$ are closed with respect to $\bm{\sigma}_{\mathrm{c}}$ [see Eqs. (67)-(69)].
We further analyze the fluctuations inherent in the estimates of the quantum Kalman filter $e_{t}=\{\braket{\hat{x}}_{\mathrm{c}},\braket{\hat{p}}_{\mathrm{c}}\}$. The variance and covariance of the estimates are defined as $\sigma_{xx,{\mathrm{m}}}\coloneqq\mathbb{E}_{e_{t}}[\braket{\hat{x}}_{\mathrm{c}}^{2}]-(\mathbb{E}_{e_{t}}[\braket{\hat{x}}_{\mathrm{c}}])^{2}$, $\sigma_{pp,{\mathrm{m}}}\coloneqq\mathbb{E}_{e_{t}}[\braket{\hat{p}}_{\mathrm{c}}^{2}]-(\mathbb{E}_{e_{t}}[\braket{\hat{p}}_{\mathrm{c}}])^{2}$, $\sigma_{xp,{\mathrm{m}}}\coloneqq\mathbb{E}_{e_{t}}[\braket{\hat{x}}_{\mathrm{c}}\braket{\hat{p}}_{\mathrm{c}}]-\mathbb{E}_{e_{t}}[\braket{\hat{x}}_{\mathrm{c}}]\mathbb{E}_{e_{t}}[\braket{\hat{p}}_{\mathrm{c}}]$. We define the covariance matrix of these as $\bm{\sigma}_{\mathrm{m}}$. The covariance matrix can be decomposed as $\bm{\sigma}=\bm{\sigma}_{\mathrm{c}}+\bm{\sigma}_{\mathrm{m}}$. From Eqs. (22) and (23), the differential equations governing the time evolution of $\bm{\sigma}_{\mathrm{m}}$ are obtained as:
$\frac{d}{dt}\sigma_{xx,{\mathrm{m}}}=$ $2\omega\sigma_{xp,{\mathrm{m}}}-2a_{x}\sigma_{xx,{\mathrm{m}}}-2a_{p}\sigma_{xp,{\mathrm{m}}}$
$-\gamma\sigma_{xx,{\mathrm{m}}}+8k\eta\sigma_{xx,\mathrm{c}}^{2}$ (24)
$\frac{d}{dt}\sigma_{pp,{\mathrm{m}}}=$ $-2\omega\sigma_{xp,{\mathrm{m}}}-2b_{x}\sigma_{xp,{\mathrm{m}}}-2b_{p}\sigma_{pp,{\mathrm{m}}}$
$-\gamma\sigma_{pp,{\mathrm{m}}}+8k\eta\sigma_{xp,{\mathrm{c}}}^{2}$ (25)
$\frac{d}{dt}\sigma_{xp,{\mathrm{m}}}=$ $-(\omega+b_{x})\sigma_{xx,{\mathrm{m}}}+(\omega-a_{p})\sigma_{pp,{\mathrm{m}}}$
$-(a_{x}+b_{p}+\gamma)\sigma_{xp,{\mathrm{m}}}+8k\eta\sigma_{xx,\mathrm{c}}\sigma_{xp,\mathrm{c}}$ (26)
In these equations, terms involving $a_{x}$, $a_{p}$, $b_{x}$, $b_{p}$ represent the contributions from feedback control. We note that the steady-state value of $\bm{\sigma}_{\rm c}$ does not depend on feedback gains [see Eqs. (67)-(69)]. Therefore, the reduction of $\braket{n}$ is solely determined by how much $\bm{\sigma}_{\mathrm{m}}$ can be reduced via feedback control. For simplicity, let $a_{x}=b_{p}=g$, $a_{p}=b_{x}=0$, and assume that $g$ is a constant. Under these conditions, it can be shown that, in the limit $g\to\infty$, each component of $\bm{\sigma}_{\mathrm{m}}$ scales as $O(1/g)$. This implies that $\bm{\sigma}_{\mathrm{m}}$ can be arbitrarily small in the limit of an infinite feedback rate. Since $\bm{\sigma}\rightarrow\bm{\sigma}_{\mathrm{c}}$ holds in this regime, the cooling limit determined by (21) is achieved. In this case, $\hat{\rho}_{t}^{e_{t}}=\hat{\rho}_{t}$ also holds and the equality in (18) is achieved. Moreover, since Eq. (20) holds in the steady state with quantum Kalman filter, the equality in (14) is also achieved.
We now prove the asymptotic property $\bm{\sigma}_{\mathrm{m}}=O(1/g)$. The time evolution equations for $\bm{\sigma}_{\mathrm{m}}$, as described by Eqs. (24), (25), and (26), can be rewritten in vector form as follows:
$\frac{d}{dt}\bm{s}_{\mathrm{m}}=\bm{A}\bm{s}_{\mathrm{m}}+\bm{v},$ (27)